CALL FOR PAPERS CALL FOR PAPERS

(revised 1 January 1997)

The Chicago Journal of Theoretical Computer Science The MIT Press

The Chicago Journal of Theoretical Computer Science is a peer-reviewed scholarly journal in theoretical computer science. The journal is committed to providing a forum for significant results on theoretical aspects of all topics in Computer Science, with a particular emphasis on the areas of research of the editorial board, listed below.

Editor in chief: Janos Simon (simon@cs.uchicago.edu). Consulting editors:

Joseph Halpern (halpern@cs.cornell.edu) Stuart A. Kurtz (kurtz@cs.uchicago.edu) Raimund Seidel (seidel@cs.uni-sb.de, seidel@cs.berkeley.edu)

Editors:

Martin Abadi  Ronald Fagin  Neil Immerman James Royer
Pankaj Agarwal Lance Fortnow  Howard Karloff  Nir Shavit
Eric Allender  Steven Fortune  Philip Klein  Alan Selman
Tetsuo Asano  Greg Frederickson Phokion Kolaitis Eva Tardos
Laszló Babai  Andrew Goldberg Stephen Mahaney Sam Toueg
Eric Bach Georg Gottlob Michael Merritt Moshe Vardi
Stephen Brookes Vassos Hadzilacos John Mitchell  Jennifer Welch
Jin-Yi Cai  Juris Hartmanis Ketan Mulmuley Pierre Wolper
Anne Condon  Maurice Herlihy Gil Neiger
Cynthia Dwork Ted Herman  David Peleg
David Eppstein Stephen Homer  Andrew Pitts

Managing editor: Michael J. O’Donnell (odonnell@cs.uchicago.edu).

Journal Format. Articles are submitted and published in L^{A}TEX source form, using the AMS-L^{A}TEX packages when appropriate, and distributed internationally over the Internet. Articles are augmented by refereed forward references to improvements and subsequent related work. Readers may obtain articles through FTP and HTTP (World Wide Web) (e.g., using Mosaic or Netscape). Other widely used network tools will be supported as they arise in the future. The Journal is committed to minimizing publication delays, and to promoting maximum flexibility in the ways that readers use the journal for teaching, research, and scholarship. Readers’ license is limited only as required to insure fair attribution to authors and the journal, and to prohibit use in a competing commercial product. In addition, the journal will publish, when appropriate, papers with unusual requirements, such as extreme length, experimental data, or computer programs.

Submission of Articles. Articles are submitted in L^{A}TEX source format, with separate BibTEX bibliography, by Internet mail to

     chicago-journal@cs.uchicago.edu

Full instructions for authors are available on the network, as described below.

Reviewing, Revision, Publication. The editors of the Journal will review each new or revised article as promptly as possible, returning one or more referee reports to the contact author and a decision to accept, reject, or require revisions. Upon acceptance, the Journal requires 4 weeks for copy editing, after which the L^{A}TEX source of the article is placed immediately in the network servers for access by readers.

More Information. More information, including detailed instructions for authors, and L^{A}TEX style files to aid authors and readers, is available through

• World Wide Web

       – http://www-mitpress.mit.edu/jrnls-catalog/science-toc.html
       – http://cs-www.uchicago.edu/publications/cjtcs

• Anonymous ftp

       – server mitpress.mit.edu, path pub/CJTCS
       – server cs.uchicago.edu, path pub/publications/cjtcs

• Network mail

       – journals-info@mit.edu
       – chicago-journal@cs.uchicago.edu